# Ceresa cycles of bielliptic Picard curves

### v1

@@ 4.7. Motives of abelian schemes
Let $A\rightarrow S$ be an abelian scheme of relative dimension $g$. Deninger–Murre [13, §3] have shown that the motive $\mathfrak{h}(A/S)$ has a canonical decomposition into Chow–Kunneth components.
@@
We call the decomposition (4.1) the canonical Chow–Kunneth decomposition of $\mathfrak{h}(A/S)$. In what follows $\mathfrak{h}^{i}(A/S)$ will always denote the $i$-th component of this decomposition.
@@
[24, Proposition 2.2.1] Let $A,B\rightarrow S$ be abelian schemes of the same relative dimension. Then the pullback map
@@
We also include the following result on abelian schemes, which follows from [17, §2, Lemma 1] and the fact that adequate schemes are normal.
@@ 5.1. An isomorphism of motives
Let $C=C_{a,b}$ be a bielliptic Picard curve (3.1) over a field $k$ with Jacobian $J$. The $\mu_{6}$-action on $C$ induces, via pullback, a $\mu_{6}$-action on $J$ and $\mathfrak{h}(J)$. The uniqueness of the Chow–Kunneth decomposition (4.1) shows that this $\mu_{6}$-action restricts to a $\mu_{6}$-action on $\mathfrak{h}^{i}(J)$ for each $0\leq i\leq 6$. Let $E^{\Delta}\colon y^{2}=x^{3}+4b(a^{2}-4b)^{2}$ be the sextic twist of $E:y^{2}=x^{3}+16(a^{2}-4b)$ by $\Delta=\Delta_{a,b}=16b(a^{2}-4b)$. Recall from §4.5 that $\mathfrak{h}^{3}(J)^{\mu_{6}}$ is a direct summand of $\mathfrak{h}^{3}(J)$.
@@
Recall the finite etale cover $\tilde{S}\rightarrow S$ from §3.3 given by adjoining a sixth root of unity and a sixth root of the discriminant $\Delta$. This cover has Galois group $G={\rm Aut}(\tilde{S}/S)\simeq D_{6}$. Also recall from Theorem 3.4 the $G$-representation ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}$ of self-dual homomorphisms; let ${\rm NS}_{\mathcal{P}}$ be the corresponding Artin motive in $\mathsf{Mot}(S)$ under the equivalence of Lemma 4.3. The $\mu_{3}$-action on $\mathcal{P}$ induces via pullback a $\mu_{3}$-action on ${\rm NS}_{\mathcal{P}}$ which is described in Lemma 3.5.
@@
By Corollary 4.2, it suffices to find a $G$-stable decomposition of $\mathfrak{h}^{2}(\mathcal{P}_{\tilde{S}}/\tilde{S})$ with properties analogous to $(1)$ and $(2)$. Theorem 3.4 shows that $B={\rm End}^{0}(\mathcal{P}_{\tilde{S}})$ is a nonsplit quaternion algebra. By functoriality we obtain a homomorphism $B^{\mathrm{op}}\rightarrow{\rm End}(\mathfrak{h}^{1}(\mathcal{P}_{\tilde{S}}/\tilde{S}))$, where $B^{\mathrm{op}}$ is the opposite algebra of $B$. We identify $B^{\mathrm{op}}$ with $B$ using the canonical involution on $B$. Moonen [32, Theorem 7.2] has shown that this $B$-action on $\mathfrak{h}^{1}(\mathcal{P}_{\tilde{S}}/\tilde{S})$ can be used to refine the Chow–Kunneth decomposition of $\mathfrak{h}(\mathcal{P}_{\tilde{S}}/\tilde{S})$. Our specific situation is explained in detail in [32, §8.2]: let
@@
be the decomposition described in [32, Bottom of p. 104]. There exist algebraic representations $\rho_{1},\rho_{2}$ of $B^{\times}$ such that ${\rm Hom}(M,R^{(2,0)})$ is $\rho_{1}$-isotypic and ${\rm Hom}(M,R^{(1,1)})$ is $\rho_{2}$-isotypic for every motive $M$ in $\mathsf{Mot}(S)$. The representation $\rho_{1}$ is a twist of the conjugation action of $B^{\times}$ on trace zero elements in $B$, and the representation $\rho_{2}$ is the norm map $\text{Nm}\colon B^{\times}\rightarrow{\rm\mathbb{Q}}^{\times}$. The $\mu_{3}$-action on $\mathcal{P}_{\tilde{S}}$ is induced from restricting the $B^{\times}$-action to a quadratic field ${\rm\mathbb{Q}}(\omega)^{\times}\subset B^{\times}$ and further restricting to its subgroup $\mu_{3}=\langle\omega\rangle\subset{\rm\mathbb{Q}}(\omega)^{\times}$. Since Nm is trivial on $\mu_{3}$, it follows that $\mu_{3}$ acts trivially on $R^{(1,1)}$. By construction of $R^{(2,0)}$ [32, §6.4], there is a canonical isomorphism $\alpha\colon\mathbb{L}\otimes({\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}_{\tilde{S}}^{\vee})\otimes{\rm\mathbb{Q}})\xrightarrow{\sim}R^{(2,0)}$. Since $G$ preserves the isomorphism classes of $\rho_{1}$ and $\rho_{2}$, the decomposition (5.3) is $G$-stable. Therefore by Corollary 4.2, $R^{(2,0)}\oplus R^{(1,1)}$ descends to a decomposition $\mathfrak{h}_{\mathrm{alg}}^{2}(\mathcal{P}/S)\oplus\mathfrak{h}^{2}_{\mathrm{tr}}(\mathcal{P}/S)$. Since the isomorphism $\alpha$ is $G$-equivariant, it descends to an isomorphism ${\rm NS}_{\mathcal{P}}(-1)\simeq\mathfrak{h}^{2}_{\mathrm{alg}}(\mathcal{P}/S)$, as desired. ∎
For the remainder of the paper we will fix a decomposition (5.4). (It is in fact unique, but we will not use this.)
@@
To analyze the piece ${\rm NS}_{\mathcal{P}}\otimes\mathfrak{h}^{1}(\mathcal{E}/S)$, we define and study an auxiliary abelian scheme $\mathcal{E}\otimes\mathrm{std}$. Recall from §3.3 the finite etale cover $\tilde{S}\rightarrow S$ with Galois group $G\simeq D_{6}$ and its two-dimensional representation $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Q}})$, which in fact lands in ${\rm GL}_{2}({\rm\mathbb{Z}})$. Let $\mathcal{E}\otimes\mathrm{std}$ be the (up to isomorphism) unique abelian scheme over $S$ that is isomorphic to $\mathcal{E}^{2}$ over $\tilde{S}$ with descent data given by $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Z}})\subset{\rm End}(\mathcal{E}^{2})$. (This is the analogue of the construction of [31] in the relative setting.)
@@ 5.2. Chow groups
Our next goal is to translate the isomorphism of Theorem 5.2 into a concrete statement about Chow groups. The (covariant) Chow groups of a motive $M$ in $\mathsf{Mot}(S)$ are by definition the ${\rm\mathbb{Q}}$-vector spaces ${\rm CH}^{p}(M):={\rm Hom}(\mathbb{L}^{\otimes p},M)$ for $p\in{\rm\mathbb{Z}}$, where we recall from §4.2 that $\mathbb{L}$ denotes the Lefchetz motive. For example, ${\rm CH}^{p}(\mathfrak{h}(X/S))={\rm CH}^{p}(X;{\rm\mathbb{Q}})$ for all $X/S$ in $\mathsf{SmProj}(S)$. We will compute ${\rm CH}^{2}$ of $\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}}$ and $\mathfrak{h}^{1}(\mathcal{E}^{\Delta}/S)(-1)$ below. To describe the former we will use the Beauville decomposition of the Chow groups of an abelian variety [3], extended by Deninger–Murre to the relative setting [13, Theorem 2.19].
@@
    We have ${\rm CH}^{2}(\mathfrak{h}^{1}(\mathcal{E}^{\Delta}_{T}/T)(-1))={\rm CH}^{1}(\mathfrak{h}^{1}(\mathcal{E}^{\Delta}_{T}/T))$. We will prove the stronger statement that if $X\rightarrow T$ is an abelian scheme of relative dimension $1$ then ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))=X(T)\otimes{\rm\mathbb{Q}}$. This is the relative version of a classical curve computation [34, Theorem 2.7.2]; we briefly sketch the details. Write $\Gamma_{e}\subset X\times_{T}X$ for the graph of the zero section $e\colon T\rightarrow X$. Then $\mathfrak{h}^{1}(X/T)=(X/T,[\Delta_{X/T}]-[\Gamma_{e}]-[^{t}\Gamma_{e}],0)$. Using the equivalence between Cartier divisors and line bundles and unwinding the definition of ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))$, we see that it is exactly the ${\rm\mathbb{Q}}$-tensor product of the set of isomorphism classes of line bundles on $X$, fiberwise of degree zero and rigidified along the zero section. In other words, ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))=\mathbf{Pic}^{0}_{X/T}(T)\otimes{\rm\mathbb{Q}}$, where $\mathbf{Pic}^{0}_{X/T}$ denotes the identity component of the Picard functor [10, §8.1, Proposition 4]. Since $X$ is a relative elliptic curve, the map $P\mapsto[P]-[e]$ induces an isomorphism $X\simeq\mathbf{Pic}^{0}_{X/T}$.
@@
    By definition of the Chow–Kunneth decomposition of Theorem 4.4 and the Beauville decomposition (5.5), ${\rm CH}^{p}(\mathfrak{h}^{i}(A/S))={\rm CH}^{p}_{(2p-i)}(A)$ for every abelian scheme $A/S$ and $i,p\in{\rm\mathbb{Z}}$. Therefore ${\rm CH}^{2}(\mathfrak{h}^{3}(\mathcal{J}_{T}/T))={\rm CH}^{2}_{(1)}(\mathcal{J}_{T};{\rm\mathbb{Q}})$. Since ${\rm CH}^{p}(M^{G})={\rm CH}^{p}(M)^{G}$ for every finite group $G$ acting on a motive $M$, the result follows from taking $\mu_{6}$-fixed points.
@@
{tikzcd}
@@ 5.3. The image of the Ceresa cycle
$\kappa(\mathcal{C}):=[\mathcal{C}]-(-1)^{*}[\mathcal{C}]\in{\rm CH}^{2}(\mathcal{J}).$
For every field-valued point $s\colon{\rm Spec}(k)\rightarrow S$, the pullback $s^{*}(\kappa(\mathcal{C}))\in{\rm CH}^{2}(\mathcal{J}_{s})$ equals the (canonical) Ceresa cycle $\kappa(\mathcal{C}_{s})$ of the curve $\mathcal{C}_{s}$ over $k$ with respect to the basepoint at infinity defined in §2.1. The next lemma shows that ${\rm CH}^{2}$ of $\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}}$ naturally contains the Ceresa cycle.
@@
Write $C=\mathcal{C}_{s}$ and $J=\mathcal{J}_{s}$. Since $J$ is defined over a field, Beauville has shown [3, Theoreme and Proposition 3(a)] that ${\rm CH}^{2}_{(1)}(J;{\rm\mathbb{Q}})$ is zero unless $i\in\{0,1,2\}$. So using (5.5) we may write $[C]=[C]_{0}+[C]_{1}+[C]_{2}$ where $(n)^{*}$ acts on $[C]_{i}$ as $n^{4-i}$. Therefore $\kappa(C)=[C]-(-1)^{*}[C]=([C]_{0}+[C]_{1}+[C]_{2})-([C]_{0}-[C]_{1}+[C]_{2})=2[C]_{1}\in{\rm CH}^{2}_{(1)}(J;{\rm\mathbb{Q}})$. Since $\kappa(C)$ is also $\mu_{6}$-invariant, we conclude the lemma. ∎
Let $\kappa(\mathcal{C})_{(1)}$ be the component in ${\rm CH}^{2}_{(1)}(\mathcal{J})$ of $\kappa(\mathcal{C})$ in the Beauville decomposition (5.5) of ${\rm CH}^{2}(\mathcal{J})$. (It might very well be that $\kappa(\mathcal{C})=\kappa(\mathcal{C})_{(1)}$, but this is not needed in our analysis.) Since $\kappa(\mathcal{C})$ is $\mu_{6}$-invariant, $\kappa(\mathcal{C})_{(1)}\in{\rm CH}^{2}_{(1)}(\mathcal{J})^{\mu_{6}}={\rm CH}^{2}(\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}})$. Choose for the remainder of the paper a collection of isomorphisms $\Phi_{T}$ satisfying the conclusions of Corollary 5.11. Define
@@
Let $\text{den}(\sigma_{0})$ be the smallest positive integer such that $\text{den}(\sigma_{0})\cdot\sigma_{0}\in\mathcal{E}^{\Delta}(S)$. Let $\sigma=\text{den}(\sigma_{0})\cdot\sigma_{0}$, a section of the abelian scheme $\mathcal{E}^{\Delta}\rightarrow S$.
@@
For every field $k$ and $s\in S(k)$, $\sigma(s)\in\mathcal{E}^{\Delta}_{s}(k)$ has infinite order if and only if $\kappa(\mathcal{C}_{s})\in{\rm CH}^{2}(\mathcal{J}_{s})$ is of infinite order.

### v2

@@ 4.8. Motives of abelian schemes
Let $A\rightarrow S$ be an abelian scheme of relative dimension $g$. Deninger–Murre [15, §3] have shown that the motive $\mathfrak{h}(A/S)$ has a canonical decomposition into Chow–Künneth components.
@@
We call the decomposition (4.1) the canonical Chow–Künneth decomposition of $\mathfrak{h}(A/S)$. In what follows $\mathfrak{h}^{i}(A/S)$ will always denote the $i$-th component of this decomposition.
@@
[28, Proposition 2.2.1] Let $A,B\rightarrow S$ be abelian schemes of the same relative dimension. Then the pullback map
@@
We also include the following result on abelian schemes, which follows from [19, §2, Lemma 1] and the fact that adequate schemes are normal.
@@ 5.1. An isomorphism of motives
Let $C=C_{a,b}$ be a bielliptic Picard curve (3.1) over a field $k$ with Jacobian $J$. The $\mu_{6}$-action on $C$ induces, via pullback, a $\mu_{6}$-action on $J$ and $\mathfrak{h}(J)$. The uniqueness of the Chow–Künneth decomposition (4.1) shows that this $\mu_{6}$-action restricts to a $\mu_{6}$-action on $\mathfrak{h}^{i}(J)$ for each $0\leq i\leq 6$. Let $E^{\Delta}\colon y^{2}=x^{3}+4b(a^{2}-4b)^{2}$ be the sextic twist of $E:y^{2}=x^{3}+16(a^{2}-4b)$ by $\Delta=\Delta_{a,b}=16b(a^{2}-4b)$. Recall from §4.5 that $\mathfrak{h}^{3}(J)^{\mu_{6}}$ is a direct summand of $\mathfrak{h}^{3}(J)$.
@@
Recall the finite étale cover $\tilde{S}\rightarrow S$ from §3.3 given by adjoining a sixth root of unity and a sixth root of the discriminant $\Delta$. This cover has Galois group $G={\rm Aut}(\tilde{S}/S)\simeq D_{6}$. Also recall from Theorem 3.4 the $G$-representation ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}$ of self-dual homomorphisms; let ${\rm NS}_{\mathcal{P}}$ be the corresponding Artin motive in $\mathsf{Mot}(S)$ under the equivalence of Lemma 4.3. The $\mu_{3}$-action on $\mathcal{P}$ induces via pullback a $\mu_{3}$-action on ${\rm NS}_{\mathcal{P}}$ which is described in Lemma 3.5.
@@
By Corollary 4.2, it suffices to find a $G$-stable decomposition of $\mathfrak{h}^{2}(\mathcal{P}_{\tilde{S}}/\tilde{S})$ with properties analogous to $(1)$ and $(2)$. Theorem 3.4 shows that $B={\rm End}^{0}(\mathcal{P}_{\tilde{S}})$ is a nonsplit quaternion algebra. By functoriality we obtain a homomorphism $B^{\mathrm{op}}\rightarrow{\rm End}(\mathfrak{h}^{1}(\mathcal{P}_{\tilde{S}}/\tilde{S}))$, where $B^{\mathrm{op}}$ is the opposite algebra of $B$. We identify $B^{\mathrm{op}}$ with $B$ using the canonical involution on $B$. Moonen [37, Theorem 7.2] has shown that this $B$-action on $\mathfrak{h}^{1}(\mathcal{P}_{\tilde{S}}/\tilde{S})$ can be used to refine the Chow–Künneth decomposition of $\mathfrak{h}(\mathcal{P}_{\tilde{S}}/\tilde{S})$. Our specific situation is explained in detail in [37, §8.2]: let
@@
be the decomposition described in [37, Bottom of p. 104]. There exist algebraic representations $\rho_{1},\rho_{2}$ of $B^{\times}$ such that ${\rm Hom}(M,R^{(2,0)})$ is $\rho_{1}$-isotypic and ${\rm Hom}(M,R^{(1,1)})$ is $\rho_{2}$-isotypic for every motive $M$ in $\mathsf{Mot}(S)$. The representation $\rho_{1}$ is a twist of the conjugation action of $B^{\times}$ on trace zero elements in $B$, and the representation $\rho_{2}$ is the norm map $\text{Nm}\colon B^{\times}\rightarrow{\rm\mathbb{Q}}^{\times}$. The $\mu_{3}$-action on $\mathcal{P}_{\tilde{S}}$ is induced from restricting the $B^{\times}$-action to a quadratic field ${\rm\mathbb{Q}}(\omega)^{\times}\subset B^{\times}$ and further restricting to its subgroup $\mu_{3}=\langle\omega\rangle\subset{\rm\mathbb{Q}}(\omega)^{\times}$. Since Nm is trivial on $\mu_{3}$, it follows that $\mu_{3}$ acts trivially on $R^{(1,1)}$. By construction of $R^{(2,0)}$ [37, §6.4], there is a canonical isomorphism $\alpha\colon\mathbb{L}\otimes({\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}_{\tilde{S}}^{\vee})\otimes{\rm\mathbb{Q}})\xrightarrow{\sim}R^{(2,0)}$. Since $G$ preserves the isomorphism classes of $\rho_{1}$ and $\rho_{2}$, the decomposition (5.3) is $G$-stable. Therefore by Corollary 4.2, $R^{(2,0)}\oplus R^{(1,1)}$ descends to a decomposition $\mathfrak{h}_{\mathrm{alg}}^{2}(\mathcal{P}/S)\oplus\mathfrak{h}^{2}_{\mathrm{tr}}(\mathcal{P}/S)$. Since the isomorphism $\alpha$ is $G$-equivariant, it descends to an isomorphism ${\rm NS}_{\mathcal{P}}(-1)\simeq\mathfrak{h}^{2}_{\mathrm{alg}}(\mathcal{P}/S)$, as desired. ∎
For the remainder of the paper we will fix a decomposition (5.2). It is in fact unique, but we will not use this.
@@
To analyze the piece ${\rm NS}_{\mathcal{P}}\otimes\mathfrak{h}^{1}(\mathcal{E}/S)$, we define and study an auxiliary abelian scheme $\mathcal{E}\otimes\mathrm{std}$. Recall from §3.3 the finite étale cover $\tilde{S}\rightarrow S$ with Galois group $G\simeq D_{6}$ and its two-dimensional representation $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Q}})$, which in fact lands in ${\rm GL}_{2}({\rm\mathbb{Z}})$. Let $\mathcal{E}\otimes\mathrm{std}$ be the (up to isomorphism) unique abelian scheme over $S$ that is isomorphic to $\mathcal{E}^{2}$ over $\tilde{S}$ with descent data given by $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Z}})\subset{\rm End}(\mathcal{E}^{2})$. (This is the analogue of the construction of [36] in the relative setting.)
@@ 5.2. Chow groups
Our next goal is to translate the isomorphism of Theorem 5.2 into a concrete statement about Chow groups. We will compute ${\rm CH}^{2}$ of $\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}}$ and $\mathfrak{h}^{1}(\mathcal{E}^{\Delta}/S)(-1)$ (in the sense of §4.6) below. To this end we will use the Beauville decomposition of the Chow groups of an abelian variety [3], extended by Deninger–Murre to the relative setting [15, Theorem 2.19].
@@
    We have ${\rm CH}^{2}(\mathfrak{h}^{1}(\mathcal{E}^{\Delta}_{T}/T)(-1))={\rm CH}^{1}(\mathfrak{h}^{1}(\mathcal{E}^{\Delta}_{T}/T))$. We will prove the stronger statement that if $X\rightarrow T$ is an abelian scheme of relative dimension $1$ then ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))=X(T)\otimes{\rm\mathbb{Q}}$. This is the relative version of a classical curve computation [39, Theorem 2.7.2]; we briefly sketch the details. Write $\Gamma_{e}\subset X\times_{T}X$ for the graph of the zero section $e\colon T\rightarrow X$. Then $\mathfrak{h}^{1}(X/T)=(X/T,[\Delta_{X/T}]-[\Gamma_{e}]-[^{t}\Gamma_{e}],0)$. Using the equivalence between Cartier divisors and line bundles and unwinding the definition of ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))$, we see that it is exactly the ${\rm\mathbb{Q}}$-tensor product of the set of isomorphism classes of line bundles on $X$, fiberwise of degree zero and rigidified along the zero section. In other words, ${\rm CH}^{1}(\mathfrak{h}^{1}(X/T))=\mathbf{Pic}^{0}_{X/T}(T)\otimes{\rm\mathbb{Q}}$, where $\mathbf{Pic}^{0}_{X/T}$ denotes the identity component of the Picard functor [10, §8.1, Proposition 4]. Since $X$ is a relative elliptic curve, the map $P\mapsto[P]-[e]$ induces an isomorphism $X\simeq\mathbf{Pic}^{0}_{X/T}$.
@@
    By definition of the Chow–Künneth decomposition of Theorem 4.4 and the Beauville decomposition (5.5), ${\rm CH}^{p}(\mathfrak{h}^{i}(A/S))={\rm CH}^{p}_{(2p-i)}(A)$ for every abelian scheme $A/S$ and $i,p\in{\rm\mathbb{Z}}$. Therefore ${\rm CH}^{2}(\mathfrak{h}^{3}(\mathcal{J}_{T}/T))={\rm CH}^{2}_{(1)}(\mathcal{J}_{T};{\rm\mathbb{Q}})$. Since ${\rm CH}^{p}(M^{G})={\rm CH}^{p}(M)^{G}$ for every finite group $G$ acting on a motive $M$ (see §4.6), the result follows from taking $\mu_{6}$-fixed points.
@@
${{\rm CH}^{2}_{(1)}(\mathcal{J}_{T};{\rm\mathbb{Q}})^{\mu_{6}}}$${\mathcal{E}^{\Delta}(T)\otimes{\rm\mathbb{Q}}}$${{\rm CH}^{2}_{(1)}(\mathcal{J}_{T^{\prime}};{\rm\mathbb{Q}})^{\mu_{6}}}$${\mathcal{E}^{\Delta}(T^{\prime})\otimes{\rm\mathbb{Q}}}$$\scriptstyle{\Phi_{T}}$$\scriptstyle{f^{*}}$$\scriptstyle{f^{*}}$$\scriptstyle{\Phi_{T^{\prime}}}$
@@ 5.3. The image of the Ceresa cycle
$\kappa_{\infty}(\mathcal{C}):=[\mathcal{C}]-(-1)^{*}[\mathcal{C}]\in{\rm CH}^{2}(\mathcal{J}).$
For every field-valued point $s\colon{\rm Spec}(k)\rightarrow S$, the pullback $s^{*}(\kappa_{\infty}(\mathcal{C}))\in{\rm CH}^{2}(\mathcal{J}_{s})$ equals the Ceresa cycle $\kappa_{\infty}(\mathcal{C}_{s})$ of the curve $\mathcal{C}_{s}$ over $k$ with respect to the basepoint at infinity defined in §2.1. We will work with ${\rm\mathbb{Q}}$-coefficients, so let $\kappa(\mathcal{C})$ be the image of $\kappa_{\infty}(\mathcal{C})$ in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$, and for each $s\in S(k)$ let $\kappa(\mathcal{C}_{s})$ be the image of $\kappa_{\infty}(\mathcal{C}_{s})$ in ${\rm CH}^{2}(\mathcal{J}_{s};{\rm\mathbb{Q}})$. The next lemma shows that, at least fiberwise, ${\rm CH}^{2}$ of $\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}}$ naturally contains the Ceresa cycle.
@@
Write $C=\mathcal{C}_{s}$ and $J=\mathcal{J}_{s}$. Since $J$ is defined over a field, Beauville has shown [3, Théorème and Proposition 3(a)] that ${\rm CH}^{2}_{(1)}(J;{\rm\mathbb{Q}})$ is zero unless $i\in\{0,1,2\}$. So using (5.5) we may write $[C]=[C]_{0}+[C]_{1}+[C]_{2}$ in ${\rm CH}^{2}(J;{\rm\mathbb{Q}})$, where $(n)^{*}$ acts on $[C]_{i}$ as $n^{4-i}$. Therefore $\kappa(C)=[C]-(-1)^{*}[C]=([C]_{0}+[C]_{1}+[C]_{2})-([C]_{0}-[C]_{1}+[C]_{2})=2[C]_{1}\in{\rm CH}^{2}_{(1)}(J;{\rm\mathbb{Q}})$. Since $\kappa(C)$ is also $\mu_{6}$-invariant, we conclude the lemma. ∎
Let $\kappa(\mathcal{C})_{(1)}$ be the component in ${\rm CH}^{2}_{(1)}(\mathcal{J})$ of $\kappa(\mathcal{C})$ in the Beauville decomposition (5.5) of ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$. (It might very well be that $\kappa(\mathcal{C})=\kappa(\mathcal{C})_{(1)}$, but this is not needed in our analysis.) Since $\kappa(\mathcal{C})$ is $\mu_{6}$-invariant, $\kappa(\mathcal{C})_{(1)}\in{\rm CH}^{2}_{(1)}(\mathcal{J})^{\mu_{6}}={\rm CH}^{2}(\mathfrak{h}^{3}(\mathcal{J}/S)^{\mu_{6}})$.
Choose for the remainder of the paper a collection of isomorphisms $\Phi_{T}$ satisfying the conclusions of Corollary 5.11. Define
@@
Let $\text{den}(\sigma_{0})$ be the smallest positive integer $d$ such that $d\cdot\sigma_{0}\in\mathcal{E}^{\Delta}(S)$. Let $\sigma=\text{den}(\sigma_{0})\cdot\sigma_{0}$, a section of the abelian scheme $\mathcal{E}^{\Delta}\rightarrow S$.
@@
For every field $k$ and $s\in S(k)$, $\sigma(s)\in\mathcal{E}^{\Delta}_{s}(k)$ is torsion if and only if $\kappa_{\infty}(\mathcal{C}_{s})\in{\rm CH}^{2}(\mathcal{J}_{s})$ is torsion.
@@
For technical reasons when relating the torsion orders of $\kappa_{\infty}(\mathcal{C}_{s})$ and $\sigma(s)$, it will be useful to have an integral lift of a multiple of $\kappa(\mathcal{C})_{(1)}$, in the following sense:
Lemma 5.14.
There exists a class $\kappa_{{\rm\mathbb{Z}}}\in{\rm CH}^{2}(\mathcal{J})$ and an integer $N\geq 1$ such that the image of $\kappa_{{\rm\mathbb{Z}}}$ in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$ is $N\cdot\kappa(\mathcal{C})_{(1)}$ and such that $s^{*}(\kappa_{{\rm\mathbb{Z}}})=N\cdot\kappa_{\infty}(\mathcal{C}_{s})$ for every field-valued point $s\colon{\rm Spec}(k)\rightarrow S$.
Proof.
Write $\kappa(\mathcal{C})=\kappa(\mathcal{C})_{(1)}+\kappa^{\prime}$ in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$, where $\kappa^{\prime}$ is the sum of the other components in the Beauville decomposition (5.5). There exist an integer $N_{1}\geq 1$ and classes $\alpha,\beta\in{\rm CH}^{2}(\mathcal{J})$ whose images in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$ are $N_{1}\kappa(\mathcal{C})_{(1)}$ and $N_{1}\kappa^{\prime}$ respectively. Since the class $N_{1}\kappa(\mathcal{C})-\alpha-\beta$ is zero in ${\rm CH}^{2}(\mathcal{J};{\rm\mathbb{Q}})$, it must be torsion in ${\rm CH}^{2}(\mathcal{J})$, so there exists an integer $N_{2}\geq 1$ such that $N_{1}N_{2}\kappa(\mathcal{C})=N_{2}\alpha+N_{2}\beta$.
We claim that there exists an integer $N_{3}\geq 1$ such that $N_{3}\cdot s^{*}(\beta)=0$ for every field-valued point $s\colon{\rm Spec}(k)\rightarrow S$. Lemma 5.12 shows that $s^{*}(\beta)$ is torsion in ${\rm CH}^{2}(\mathcal{J}_{s})$ for every such $s$. So if $s=\eta$ is the generic point of $S$, there exists an $M\geq 1$ such that $M\cdot\eta^{*}(\beta)=0$. By spreading out, there exists an open $U\subset S$ such that $M\cdot s^{*}(\beta)=0$ for all $s\in U(k)$. Repeatedly applying this argument to the generic points of the irreducible components of $S\setminus U$ proves the claim.
The lemma follows by taking $\kappa_{{\rm\mathbb{Z}}}=N_{2}N_{3}\alpha$ and $N=N_{1}N_{2}N_{3}$. ∎